# Back to Basics: Revisiting REINFORCE Style Optimization for Learning from Human Feedback in LLMs

### 22 Feb 2024 v1

@@ 1 Introduction
Despite being the focus of considerable research effort (ouyang2022LLMRLHF; bai2022constitutional; lee2023rlaif; tunstall2023zephyr), there is a lack of consensus regarding the optimal approach to achieve this goal. Reinforcement Learning from Human Feedback (RLHF), one of the most widely regarded alignment approaches, directly borrows from traditional RL literature and uses techniques such as Proximal Policy Optimization (PPO) to maximize the reward score produced by a reward model that is typically trained as a binary classifier on pairs of completions labeled by human annotators. While PPO has become a canonical approach cemented in popularity through its usage in the seminal literature on RLHF (LearningToSummarizeHF; nakano2022webgpt; bai2022constitutional), getting PPO to work in practice is non-trivial for non-RL specialists and comes with known issues:
@@
The core question we seek to explore in this work is can we avoid the computational and optimization complexity of PPO while preserving performance? Recent works propose “RL-free” methods such as DPO (rafailov2023DPO), IPO (azar2023IPO) or iterative fine-tuning approaches to LLM preference training (yuan2023rrhf; zhao2023slichf; dong2023raft)
All these approaches attempt to answer this question by stripping all RL components from the RLHF objective and the difficulties that come with it (rafailov2023DPO; zhao2023slichf; yuan2023rrhf; liu2023statistical; dong2023raft; azar2023IPO). Iterative finetuning techniques rely solely on a powerful reward model to identify a subset of samples to train on, while DPO avoids both reinforcement learning and a seperate reward model by modifying the loss.
In contrast to these approaches, we remain in the RL paradigm, but instead return to basics. We isolate several key differences between traditional Deep-RL settings which originally motivated PPO and typical human-preference learning settings for LLMs. We note that PPO, as an approach, emphasizes stability across iterations, aiming to train an effective policy with the premise of small, stable updates. PPO was designed for a regime where off-policy gradient updates are large enough to introduce instability. This regime dominates traditional Deep-RL benchmarks (engstrom2020implementation; schulman2017proximal). However, in this work, we posit that the setting of RLHF, which involves fine-tuning a pre-trained LLM, is lacking in these characteristics.
@@
Furthermore, we revisit how learning from human preferences is formulated in the context of RL where generating each token is modeled as an action, and each partial sequence, starting with the prompt, is seen as a state. In practice, this modeling assumption for PPO method is often voided. We argue and show that the modeling of partial sequences is unnecessary since rewards from humans are only attributed to full generations, with no true rewards for any intermediary token generated. Thus, it is more appropriate and efficient to model the entire generation as a single action with the initial state determined by the prompt.
Given these observations, while keeping simplicity as a guiding principle, we explore the use of the REINFORCE estimator (williams1992simple) and its multi-sample extension REINFORCE Leave-One-Out (RLOO) (Kool2019Buy4R) to optimize the sequence-level objective. We break apart PPO and show that the most basic policy gradient algorithm, Vanilla Policy Gradient REINFORCE consistently outperforms PPO. PPO is unnecessarily complicated for a pre-trained LLM environment. Unlike PPO, we can use REINFORCE to directly optimize the full trajectory (sequence) coupled with unbiased baselines, whereas actor-critic algorithms (sutton2000policygrad), such as PPO, bootstrap off intermediary state value-functions to reduce variance at the cost of introducing bias into the estimator.
@@
    PPO is not the right tool for doing RL in RLHF. We break apart PPO and show that the most “basic” policy gradient algorithm, Vanilla Policy Gradient REINFORCE (sutton2020reinforcement), is consistently outperforming PPO by 3.2% to 20.3% interms of winrate, across all dataset and base model pairing .
@@ 2.3 REINFORCE Leave-One-Out (RLOO)
$\begin{split}\frac{1}{k}&\sum_{i=1}^{k}[R(y_{(i)},x)-\frac{1}{k-1}\sum_{j\neq{k}}R(y_{(j)},x)]\nabla\log\pi(y_{(i)}|x)\text{ for }y_{(1)},...,y_{(k)}\overset{i.i.d}{\sim}\pi_{\theta}(.|x)\\ \end{split}$
Where $k$ refers to the number of online samples generated, $\textsc{RLOO}_{k}$ considers each $y_{(i)}$ individually and uses the remaining $k-1$ samples to create an unbiased estimate of the expected return for the prompt, akin to a parameter-free value-function, but estimated at each training step. This is a much more effective baseline (as our experiments will show) than $b_{MA}$ since it’s created on-the-fly for each sample and at each training step, but comes at a cost of increased sampling time during training.
@@ 3.1 Revisiting the need for low-variance estimators
In contrast, REINFORCE uses an unbiased Monte-Carlo estimator of the trajectory return that can have high variance in theory, especially if it is approximated with only a single sample, which is not frequently preferred in traditional Deep-RL environments. Recent work has offered a plethora of evidence that REINFORCE suffers from high variance and fails in the presence of large action spaces like NLP (ranzato2015sequence; bahdanau2017actorcritic; ding2017cold; ammanabrolu2020graph; ammanabrolu-etal-2022-aligning; martin-etal-2022-learning). However, we note that these findings were based on scenarios with poor conditioning when training from random or weak initialization as opposed to warm-starting it from a strong pre-trained model.
@@ 4.1 Training Details
Datasets   We report results on the TL;DR Summarize (LearningToSummarizeHF) and Anthropic Helpful and Harmless Dialogue (bai2022AnthropicHH) datasets. The trainig split of TL;DR Summarize  [^1] https://github.com/openai/summarize-from-feedback dataset contains 116k human-written instructions and 93k human-annotated preference pairs. The preprocessed Anthropic-HH [^2] https://huggingface.co/datasets/Dahoas/full-hh-rlhf dataset contains 112k many training preference pairs.
Models   For both datasets, we use Pythia-6.9B (biderman2023pythia) as the pretrained base-model. To ablate the effect of the pre-trained model quality on the human preference optimization, we also experiment with Llama-7B (touvron2023llama) coupled with the Anthropic-HH dataset.
@@ 4.2 Evaluation
Optimization Quality   For all online methods (all methods except DPO), to measure how well the method optimizes the intrinsic objective , we report average rewards (using the training RM) on 1000 samples from the test set. To measure how well each method optimizes the extrinsic objective of aligning the models to human preference, we report simulated win rates in accordance with Alpacafarm framework (dubois2024alpacafarm) where we use GPT-4 as a proxy for human evaluation. At evaluation, we use greedy sampling unless otherwise noted.
Alignment Tax   RLHF fine-tuning is often associated with a drop in diversity and language fluency which is referred to as alignment tax (askell2021general; kirk2024understandingRLHF). Hence, we also report metrics that serve as proxies for fluency and diversity similar to (dong2023raft). To measure fluency, we report perplexity measured using the preferred completions from the test set, similar to dong2023raft. Finally, we measure the diversity coupled with length by using the average n-gram diversity (li-etal-2016-diversity).
@@ 5.1 Reward Optimization
The objective of RLOO, REINFORCE with a (moving average) baseline, RAFT, PPO, and vanilla PG is to maximize the reward score, hence we compare the success of optimization for each method. On each dataset, we use the same reward model for all the methods, thus their test reward scores are directly comparable.
@@
Table 1: Final win rates on generations for held-out test prompts in the Anthropic-HH and TL;DR Summarize datasets. Metrics are reported for the checkpoint with the highest test reward.
5.2 Simulated Win Rates
Table 5.1 presents the win rates against the original completions in TL;DR Summarize and Antrophic-HH for each method. Here, we also include DPO.
Modeling partial completions is not necessary Recall that the key distinguishing factor between Vanilla PG and REINFORCE as referred to in this work, is that while Vanilla PG treats each token as an action, REINFORCE operates on the entire generation. As seen in Table 5.1, REINFORCE with baseline is on par with Vanilla PG on both TL;DR (70.7 vs 70.4) and HH (37.9 vs 36.4) datasets when using Pythia-based model. Moreover, REINFORCE with baseline outperforms Vanilla PG in HH dataset with Llama-based model, achieving higher win rates (55.3 vs 52.3).
@@ 5.2 Simulated Win Rates
Win rates are inline with test reward scores RLOO with $k=4$ achieves the highest results, outperforming PPO by 10.3, 14.5, and 32.1 increase in win rates against the original completions for TL;DR, HH (Pythia) and HH (Llama) respectively. As the only exception, RLOO achieves the highest win rate with $k=2$ in HH dataset.
RLOO is more sample efficient than RAFT Comparing RLOO with RAFT, under the same sampling budget $k$, RLOO consistently outperforms RAFT in all datasets and models. On average, RLOO achieves win rates of 61.3 and 61.9 for $k=2$ and $k=4$, respectively, across TL;DR, HH (Pythia), and HH (Llama), while RAFT scores 56.1 and 59.5, respectively. Notably, RLOO exhibits the highest increase in win rate compared to RAFT, reaching 9.9, in HH dataset with $k=2$ and the Pythia-based model as shown in Table 5.1.

### 26 Feb 2024 v2

@@ 1 Introduction
Despite being the focus of considerable research effort (ouyang2022LLMRLHF; bai2022constitutional; lee2023rlaif; tunstall2023zephyr; khalifa2021distributional), there is a lack of consensus regarding the optimal approach to achieve this goal. Reinforcement Learning from Human Feedback (RLHF), one of the most widely regarded alignment approaches, directly borrows from traditional RL literature and uses techniques such as Proximal Policy Optimization (PPO) to maximize the reward score produced by a reward model that is typically trained as a binary classifier on pairs of completions labeled by human annotators. While PPO has become a canonical approach cemented in popularity through its usage in the seminal literature on RLHF (LearningToSummarizeHF; nakano2022webgpt; bai2022constitutional), getting PPO to work in practice is non-trivial for non-RL specialists and comes with known issues:
@@
Recent works propose “RL-free” methods such as DPO (rafailov2023DPO), IPO (azar2023IPO) or iterative fine-tuning approaches to LLM preference training (yuan2023rrhf; zhao2023slichf; dong2023raft). However, these works fail to question whether a simpler solution within an RL paradigm exists. Instead, all these approaches attempt to answer this question by stripping all RL components from RLHF and the difficulties that come with it (rafailov2023DPO; zhao2023slichf; yuan2023rrhf; liu2023statistical; dong2023raft; azar2023IPO). Iterative fine-tuning techniques rely solely on a powerful reward model to identify a subset of samples to train on, while DPO and IPO avoid both reinforcement learning and training a separate reward model by directly learning from human feedback.
In contrast to these approaches, we remain in the RL paradigm, but instead return to basics. The core question we seek to explore in this work is can we avoid the computational and optimization complexity of PPO while preserving performance?. We isolate several key differences between traditional Deep-RL settings which originally motivated PPO and typical human-preference learning settings for LLMs. We note that PPO, as an approach, emphasizes stability across iterations, aiming to train an effective policy with the premise of small, stable updates. PPO was designed for a regime where off-policy gradient updates are large enough to introduce instability. This regime dominates traditional Deep-RL benchmarks (engstrom2020implementation; schulman2017proximal). However, in this work, we posit that the setting of RLHF, which involves fine-tuning a pre-trained LLM, is lacking in these characteristics.
@@
Furthermore, we revisit how learning from human preferences is formulated in the context of RL where generating each token is modeled as an action, and each partial sequence, starting with the prompt, is seen as a state. In practice, this modeling assumption for PPO method is often voided. We argue and show that the modeling of partial sequences is unnecessary in this setting where rewards are only attributed to full generations, with no true rewards for any intermediary tokens in the generation. Thus, it is more appropriate and efficient to model the entire generation as a single action with the initial state determined by the prompt.
Given these observations, while keeping simplicity as a guiding principle, we explore the use of the REINFORCE estimator (williams1992simple) and its multi-sample extension REINFORCE Leave-One-Out (RLOO) (Kool2019Buy4R) to optimize the sequence-level objective. We break apart PPO and show that the most basic policy gradient algorithm, Vanilla Policy Gradient REINFORCE consistently outperforms PPO. PPO is unnecessarily complicated for a pre-trained LLM environment. Unlike PPO, we can use REINFORCE to directly optimize the full trajectory (sequence) return coupled with unbiased baselines, whereas actor-critic algorithms (sutton2000policygrad), such as PPO, bootstrap off intermediary state value-functions to reduce variance at the cost of introducing bias into the estimator.
@@
    PPO is not the right tool for doing RL in RLHF. We break apart PPO and show that the most “basic” policy gradient algorithm, Vanilla Policy Gradient REINFORCE (sutton2020reinforcement), is consistently outperforming PPO by 3.2% to 20.3% interms of win-rate, across all dataset and base model pairing .
@@ 2.3 REINFORCE Leave-One-Out (RLOO)
$\begin{split}\frac{1}{k}&\sum_{i=1}^{k}[R(y_{(i)},x)-\frac{1}{k-1}\sum_{j\neq{i}}R(y_{(j)},x)]\nabla\log\pi(y_{(i)}|x)\text{ for }y_{(1)},...,y_{(k)}\overset{i.i.d}{\sim}\pi_{\theta}(.|x)\\ \end{split}$
Where $k$ refers to the number of online samples generated, $\textsc{RLOO}_{k}$ considers each $y_{(i)}$ individually and uses the remaining $k-1$ samples to create an unbiased estimate of the expected return for the prompt, akin to a parameter-free value-function, but estimated at each training step. This is a much more effective baseline (as our experiments will show) than $b_{MA}$ since it’s created on-the-fly for each sample and at each training step, but comes at a cost of increased sampling time during training. We note that generating extra samples as a means of variance reduction has been proposed by concurrent work (li2023remax), but we focus on RLOO here because of the efficiency benefits of fully utilizing all samples.
@@ 3.1 Revisiting the need for low-variance estimators
In contrast, REINFORCE uses an unbiased Monte-Carlo estimator of the trajectory return that can have high variance in theory, especially if it is approximated with only a single sample, which is not frequently preferred in traditional Deep-RL environments. Recent work has offered a plethora of evidence that REINFORCE suffers from high variance and fails in the presence of large action spaces like NLP (ranzato2015sequence; bahdanau2017actorcritic; ding2017cold; ammanabrolu2020graph; ammanabrolu-etal-2022-aligning; martin-etal-2022-learning; NEURIPS2022_67496dfa). However, we note that these findings were based on scenarios with poor conditioning when training from random or weak initialization as opposed to warm-starting it from a strong pre-trained model.
@@ 4.1 Training Details
Datasets   We report results on the TL;DR Summarize (LearningToSummarizeHF) and Anthropic Helpful and Harmless Dialogue (bai2022AnthropicHH) datasets. The trainig split of TL;DR Summarize  [^1] https://github.com/openai/summarize-from-feedback dataset contains 116k human-written instructions and 93k human-annotated preference pairs. The preprocessed Anthropic-HH [^2] https://huggingface.co/datasets/Dahoas/full-hh-rlhf dataset contains 112k training preference pairs.
Models   For both datasets, we use Pythia-6.9B (biderman2023pythia) as the pretrained base-model. To ablate the effect of the pre-trained model quality on learning from human preferences, we also experiment with Llama-7B (touvron2023llama) coupled with the Anthropic-HH dataset.
@@ 4.2 Evaluation
Optimization Quality   For all online methods (all methods except DPO), to measure how well the method optimizes the intrinsic objective , we report average rewards (using the training RM) on 1000 samples from the test set. To measure how well each method optimizes the extrinsic objective of aligning the models to human preference, on the same test samples, we report simulated win-rates in accordance with Alpacafarm framework (dubois2024alpacafarm) where we use GPT-4 as a proxy for human evaluation. We measure win-rates against reference SFT completions for the TL;DR dataset, and preferred completions for the HH dataset. At evaluation, we use greedy sampling unless otherwise noted.
Alignment Tax   RLHF fine-tuning is often associated with a drop in diversity and language fluency which is referred to as alignment tax (askell2021general; kirk2024understandingRLHF). Hence, we also report metrics that serve as proxies for fluency and diversity similar to (dong2023raft). To measure fluency, we report perplexity measured using the preferred completions from the test set, similar to dong2023raft. Finally, we measure average completion length and diversity using average n-gram diversity (li-etal-2016-diversity).
@@ 5.1 Reward Optimization
The objective of RLOO, REINFORCE with a (moving average) baseline, RAFT, PPO, and vanilla PG is to maximize the reward score, hence we compare the success of optimization for each method. On each dataset and base-model pair, we use the same reward model for all the methods, thus their test reward scores are directly comparable.
@@
Table 1: Final win-rates on generations for held-out test prompts in the Anthropic-HH and TL;DR Summarize datasets. Metrics are reported for the checkpoint with the highest test reward.
5.2 Simulated win-rates
Table 5.1 presents the win-rates against the original completions in TL;DR Summarize and Antrophic-HH for each method. Here, we also include DPO.
Modeling partial completions is not necessary Recall that the key distinguishing factor between Vanilla PG and REINFORCE as referred to in this work, is that while Vanilla PG treats each token as an action, REINFORCE operates on the entire generation. As seen in Table 5.1, REINFORCE with baseline is on par with Vanilla PG on both TL;DR (70.7 vs 70.4) and HH (37.9 vs 36.4) datasets when using Pythia-based models. Moreover, REINFORCE with baseline outperforms Vanilla PG in HH dataset with a Llama-based model, achieving a higher win-rate (55.3 vs 52.3).
@@ 5.2 Simulated win-rates
Win-rates are inline with test reward scores RLOO with $k=4$ achieves the highest win-rates, outperforming PPO by 10.3, 14.5, and 32.1 for TL;DR, HH (Pythia) and HH (Llama), respectively. As the only exception, RLOO achieves the highest win-rate with $k=2$ in HH dataset.
RLOO is more sample efficient than RAFT Comparing RLOO with RAFT, under the same sampling budget $k$, RLOO consistently outperforms RAFT in all datasets and models. When averaged across the three dataset and model pairings, RLOO achieves win-rates of 61.3 and 61.9 for $k=2$ and $k=4$, respectively, while RAFT scores 56.1 and 59.5, respectively. Notably, RLOO exhibits the highest increase in win-rate compared to RAFT, up to 9.9 as in the HH dataset with $k=2$ and Pythia-based models (second column in Table 5.1).